# Double-coset zeta functions for groups acting on trees

### v1

@@ 5.2. The property ($\ast_{k}$)
Let $k\geq 1$. A group action on a tree $(G,T)$ has property $(\ast_{k})$ if, for every geodesic $(e_{1},\ldots,e_{l+k})$ in $T$ with $l\geq 1$,
@@
One checks that property ($\ast_{k}$) implies property ($\ast_{k+1}$), for every $k\geq 1$.
Remark 5.0.1.
Let $T$ be a tree and $G\leq\mathrm{Aut}(T)$ be a subgroup with the subspace topology induced by $\mathrm{Aut}(T)$. If $(G,T)$ has property ($\ast_{k}$) for some $k\geq 1$, then $G$ is non-discrete.
In detail, assume that $(G,T)$ has property $(\ast_{k})$. If we prove that all vertex-stabilisers in $G$ are infinite, then [2, Lemma 2.1] yields the claim. Let $v\in VT$. Since $T$ has no leaves, there is a ray $(e_{i})_{i\in\mathbb{Z}_{\geq 1}}$ in $T$ with $o(e_{1})=v$. For every $h\in\mathbb{Z}_{\geq 1}$, set $\mathfrak{p}_{h}=(e_{i})_{1\leq i\leq hk}$ and note that $|G_{\mathfrak{p}_{h}}\cdot(e_{hk+1},\ldots,e_{hk+k})|=|G_{\mathfrak{p}_{h}}:G_{\mathfrak{p}_{h+1}}|\geq 2$. Therefore,
$|G_{v}:G_{\mathfrak{p}_{h}}|=|G_{v}:G_{\mathfrak{p}_{1}}|\cdot\prod_{i=1}^{h-1}|G_{\mathfrak{p}_{i}}:G_{\mathfrak{p}_{i+1}}|\geq 2^{h-1}$
for every $h\geq 1$. Thus, $G_{v}$ is infinite.
@@
Let $(G,T)$ be a weakly locally $\infty$-transitive action on a locally finite tree. Denote by $\omega$ the standard edge weight on $\Gamma=G\backslash T$, consider $N_{\mathrm{edg}}=N_{\mathrm{edg}}^{\omega}$ as in Definition 4.4, and let $k\geq 1$. Then $(G,T)$ has property ($\ast_{k}$) if, and only if, $N_{\mathrm{edg}}(\rho)\geq 2$ for every path $\rho$ in $\Gamma$ of length $k+1$ which can be lifted to a geodesic in $T$.
@@
Let $k\geq 1$ and $\mathfrak{p}=(e_{1},\ldots,e_{l+k})$ be a geodesic in $T$ with $l\geq 1$. Denote by $\pi\colon T\to\Gamma$ the quotient map of $(G,T)$ and set $\pi(e_{i})=a_{i}$ for every $1\leq i\leq l+k$. By Remark 3.9.1(ii), $G_{(e_{1},\ldots,e_{l})}$ acts transitively on $\{\mathfrak{q}=(f_{1},\ldots,f_{l+k})\in\mathrm{Geod}_{T}(e_{1}\to T)\,:\,\pi(\mathfrak{q})=\pi(\mathfrak{p})\text{ and }\forall\,i\leq l,\,f_{i}=e_{i}\}$. Hence, by Remark 4.5.1,
(5.8) $\begin{array}[]{rl}&|G_{(e_{1},\ldots,e_{l})}\cdot(e_{l+1},\ldots,e_{l+k})|=|G_{(e_{1},\ldots,e_{l})}\cdot\mathfrak{p}|=\\ =&\displaystyle{\Big{|}\{\mathfrak{q}=(f_{1},\ldots,f_{l+k})\in\mathrm{Geod}_{T}(e_{1}\to T)\,:\,\pi(\mathfrak{q})=\pi(\mathfrak{p}),\forall\,i\leq l,\,f_{i}=e_{i}\}\Big{|}}\\ =&\displaystyle{\prod_{i=l}^{l+k-1}N_{\mathrm{edg}}(a_{i},a_{i+1})=N_{\mathrm{edg}}(a_{l},\ldots,a_{l+k}).}\end{array}$
This yields the “if” part of the statement. For the “only if” part, let $\rho=(a_{1},\ldots,a_{k+1})$ be an arbitrary path in $\Gamma$ which can be lifted to a geodesic $\mathfrak{p}=(e_{1},\ldots,e_{k+1})$ in $T$. Remark 4.5.1 now yields $|G_{e_{1}}\cdot\mathfrak{p}|=N_{\mathrm{edg}}(\rho)\geq 2$. ∎
@@
Let $(G,T)$ be a weakly locally $\infty$-transitive group action on a tree with quotient graph $\Gamma$ and standard edge weight $\omega$. Assume that $\omega(E\Gamma)\subseteq\mathbb{Z}_{\geq 2}$. Then the following are equivalent:
@@
Moreover $(G,T)$ has property $(\ast_{1})$ if, and only if, $\omega(E\Gamma)\subseteq\mathbb{Z}_{\geq 3}$.
@@
For the first part of the statement, by Proposition 5.1 it suffices to prove that, given an arbitrary $k\geq 2$, condition (iii) is equivalent to have $N_{\mathrm{edg}}(\rho)\geq 2$ for all paths $\rho$ in $\Gamma$ of length $k+1$. But, given a path $(a_{1},\ldots,a_{k+1})$ in $\Gamma$ with $k\geq 2$, we have $N_{\mathrm{edg}}(a_{1},\ldots,a_{k+1})=1$ if, and only if, $a_{i+1}=\bar{a}_{i}$ and $\omega(a_{i+1})=2$ for every $1\leq i\leq k$.
For the second part of the statement, one proceeds analogously. Namely, it suffices to prove that $(G,T)$ does not have property $(\ast_{1})$ if, and only if, there is a length-$2$ path $(a_{1},a_{2})$ in $\Gamma$ such that $N_{\mathrm{edg}}(a_{1},a_{2})=1$. In turn, $N_{\mathrm{edg}}(a_{1},a_{2})=1$ is equivalent to have $a_{2}=\bar{a}_{1}$ and $\omega(a_{2})=2$. ∎
@@
Let $k\geq 1$ and consider a geodesic $\mathfrak{p}=(e_{1},\ldots,e_{l+k})$ in $T$ with $l\geq 1$. By Proposition 3.1 and Corollary 4.13,
$\begin{split}|G_{(e_{1},\ldots,e_{l})}\cdot(e_{l+1},\ldots,e_{l+k})|&=|G_{e_{l}}\cdot(e_{l+1},\ldots,e_{l+k})|\\ &=\mathcal{W}(g\cdot(e_{l},\ldots,e_{l+k})),\end{split}$
for some $g\in G_{o(e_{l})}$. This yields the “if” part of the statement. For the “only if” part, let $\xi=(x_{1},\ldots,x_{k+1})\in\mathcal{P}_{(\Delta,\iota)}$ and $e\in ET^{+}$ such that $\mathcal{L}(e)=x_{1}$. By Lemma 3.5, there is a geodesic $\mathfrak{p}=(e_{1},\ldots,e_{k+1})$ in $T$ with $e_{1}=e$ such that $\mathcal{L}(\mathfrak{p})=\xi$. By Proposition 4.11 we conclude that
$|G_{e_{1}}\cdot(e_{2},\ldots,e_{k})|=\mathcal{W}(\xi)\geq 2.\qed$
@@ 5.3. Convergence properties
The main goal of what follows is to study the double-coset property and the polynomial double-coset growth of triples $(G,G_{t_{1}},G_{t_{2}})$, where $G$ is a group acting on a tree $T$ and $t_{1},t_{2}\in T$.
@@
Let $(G,T)$ be a group action on a tree.
@@
    Assume that $C:=\sup_{e,f\in ET\,:\,t(e)=o(f)}|G_{e}\cdot f|$ is finite. Then, for every geodesic $\mathfrak{p}=(e_{1},\ldots,e_{l})$ in $T$ of length $l\geq 1$, we have $|G_{e_{1}}:G_{\mathfrak{p}}|\leq C^{l-1}$.
@@
    Suppose that $(G,T)$ have property ($\ast_{k}$) for some $k\geq 1$. Then, for every geodesic $\mathfrak{p}=(e_{1},\ldots,e_{l})$ in $T$ of length $l\geq 1$, we have $|G_{e_{1}}:G_{\mathfrak{p}}|\geq 2^{\frac{l-k}{k}}$.
@@
Since $|G_{(e_{1},\ldots,e_{i})}\cdot e_{i+1}|\leq|G_{e_{i}}\cdot e_{i+1}|$ for every $11\leq i\leq l-$ 1, we obtain (i).
@@
To prove the latter inequality in (5.10), set $A_{j}=|G_{(e_{1},\ldots,e_{j})}\cdot e_{j+1}|$ for every $11\leq j\leq l-$ 1. Then the product on the right-hand side of (5.10) becomes
@@
where, for every $11\leq j\leq l-$ 1, we put
$\alpha_{j}:=|\{i\,:\,\max\{1,j-k+1\}\leq i\leq\min\{j,l-k\}\}|.$
It remains to show that $\alpha_{j}\leq k$ for every $11\leq j\leq l-$ 1. If $j\leq k-1$, then $\alpha_{j}\leq j<k$. If $k\leq j\leq l-k$, then $\alpha_{j}=|\{i\,:\,j-k+1\leq i\leq j\}|=k$. Finally, if $j\geq k$ and $j\geq l-k$ then $\alpha_{j}=|\{i\,:\,j-k+1\leq j\leq l-k\}|=l-j\leq k$. Hence (5.10) holds. Combining (5.10), the orbit-stabiliser theorem and the fact that $(G,T)$ has property $(\ast_{k})$, we conclude that
@@
Let $(G,T)$ be a group action on a locally finite tree with finite quotient graph, and set $M:=\sup_{v\in VT}|o^{-1}(v)|$. If $(G,T)$ has property $(\ast_{k})$ for some $k\geq 1$, then $3\leq M<\infty$ and, for all $t_{1},t_{2}\in T$ and $n\geq 1$,
@@
In particular, for all $t_{1},t_{2}\in~{}T$ the triple $(G,G_{t_{1}},G_{t_{2}})$ has polynomial double-coset growth.
@@
Let $t_{1},t_{2}\in T$ and consider $T_{1}$ and $T_{2}$ as defined before (5.12). We first observe that $3\leq M\leq\infty$. Clearly, $|o^{-1}(v)|=|o^{-1}(g\cdot v)|$ for all $v\in VT$ and $g\in G$. Since $T$ is locally finite and $G\backslash VT$ is finite, we have $M<\infty$. Moreover, $M\geq 2$ because $T$ is assumed to have no leaves (cf. Section 3.1). Since $(G,T)$ has property $(\ast_{k})$, then $T$ cannot be a bi-infinite line and $M\geq 3$.
@@
$\begin{split}b_{n}(G,G_{t_{1}},G_{t_{2}})&=|\{\mathfrak{p}\in\mathrm{Geod}_{T}(T_{1}\to T)\,:\,|G_{t_{1}}:G_{\mathfrak{p}}|=n\}|\\ &\leq|\{\mathfrak{p}\in\mathrm{Geod}_{T}(T_{1}\to T)\,:\,\ell(\mathfrak{p})\leq\lfloor k\cdot\log_{2}n\rfloor+k\}|\\ &=\sum_{t\in T_{1}}|\{\mathfrak{p}\in\mathrm{Geod}_{T}(t\to T)\,:\,\ell(\mathfrak{p})\leq\lfloor k\cdot\log_{2}n\rfloor+k\}|\\ &\leq\sum_{t\in T_{1}}\Bigg{(}1+\sum_{l=1}^{\lfloor k\cdot\log_{2}n\rfloor+k}M(M-1)^{l-1}\Bigg{)}\\ &=|T_{1}|\cdot\frac{M(M-1)^{\lfloor k\cdot\log_{2}n\rfloor+k}-2}{M-2}.\end{split}$
@@
Let $(G,T)$ be a weakly locally $\infty$-transitive group action on a tree with finite quotient graph $\Gamma$. Assume that the standard edge weight $\omega$ on $\Gamma$ takes values in $\mathbb{Z}_{\geq 2}$. Then the following are equivalent for all $t_{1},t_{2}\in T$:
@@
    $\omega(a)\geq 3$ or $\omega(\bar{a})\geq 3$ for every $a\in E\Gamma$.
@@
Clearly, (ii)$\Rightarrow$(i). Moreover, Proposition 5.5 and Corollary 5.2 imply (iii)$\Rightarrow$(ii) and (iii)$\Leftrightarrow$(iv), respectively. It remains to prove (i)$\Rightarrow$(iv).
Assume that there is $a\in E\Gamma$ such that $\omega(a)=\omega(\bar{a})=2$, and let $\mathfrak{p}=(a,\bar{a})$. Then $N_{\mathrm{edg}}(\mathfrak{p}^{d})=(\omega(\bar{a})-1)^{d}(\omega(a)-1)^{d-1}=1$ for every $d\geq 1$. Consider two paths $\mathfrak{q}_{1}=(a_{1},\ldots,a_{h})$ and $\mathfrak{q}_{2}=(b_{1},\ldots,b_{k})$ of positive length in $\Gamma$ from $\pi(t_{1})$ to $o(a)$ and from $o(a)$ to $\pi(t_{2})$, respectively. Then, for every $d\geq 1$ we have
@@
Let $(G,T)$ be a (P)-closed action on a locally finite tree. Assume that the quotient graph is finite and its standard edge weight takes values in $\mathbb{Z}_{\geq 2}$. Then the following are equivalent for all $t_{1},t_{2}\in T$:
@@
    $(G,T)$ has property $(\ast_{k})$ for some $k\geq 1$.
@@
The implication (ii)$\Rightarrow$(i) is immediate, and (iii)$\Rightarrow$(ii) follows from Proposition 5.5. It remains to prove (i)$\Rightarrow$(iii). Assume that $(G,T)$ does not have property $(\ast_{k})$ for every $k\geq 1$. Let $\Delta$ be the local action diagram associated to $(G,T)$ and consider an inversion $\iota$ in $\Delta$. Without loss of generality, $G=U(\Delta,\iota,c_{0})$ and $T=T(\Delta,\iota,c_{0})$ for some $c_{0}\in V\Gamma$ (cf. Theorem 3.8). By Proposition 5.3, there is a reduced path $(x_{1},\ldots,x_{k})$ in $(\Delta,\iota)$ of length $k\geq|X|^{2}+2$ such that $\mathcal{W}(x_{1},\ldots,x_{k})=1$, i.e., $\mathcal{W}(x_{i},x_{i+1})=1$ for every $11\leq i\leq k-$ 1. Since $k\geq|X|^{2}+2$, we have $(x_{i},x_{i+1})=(x_{i+l},x_{i+1+l})$ for some $i,l\geq 1$. Set $\eta:=(x_{j})_{i\leq j\leq i+l-1}$. Hence, for every $d\geq 1$ the power $\eta^{d}$ is a reduced path in $(\Delta,\iota)$ satisfying
@@
Set $y\in X_{\pi(t_{1})}$, $z\in X_{\pi(t_{2})}$ and choose arbitrary reduced paths of positive length in $(\Delta,\iota)$, say $\xi_{1}=(y_{1},\ldots,y_{h})$ and $\xi_{2}=(z_{1},\ldots,z_{r})$, such that $\xi_{1}\cdot x_{1}\in\mathcal{P}_{(\Delta,\iota)}(y\to x_{1})$ and $x_{i+l-1}\cdot\xi_{2}\in\mathcal{P}_{(\Delta,\iota)}(x_{i+l-1}\to z)$. Such reduced paths exist because the standard edge weight of $G\backslash T$ takes values in $\mathbb{Z}_{\geq 2}$ (cf. Remark 4.5.1). For every $d\geq 1$, the path $\xi_{1}\cdot\eta^{d}\cdot\xi_{2}$ is reduced in $(\Delta,\iota)$ and
@@
By Lemma 3.5, for every $d\geq 1$ there is $\tilde{\mathfrak{q}}_{d}\in\mathrm{Geod}_{T}(T_{1}\to G\cdot T_{2})$ such that $\mathcal{L}(\tilde{\mathfrak{q}}_{d})=\xi_{1}\cdot\eta^{d}\cdot\xi_{2}$. By Corollary 4.13, we may assume that all edges of $\tilde{\mathfrak{q}}_{d}$ are in $ET^{+}$ for every $n\geq 1$. Note that each $\tilde{\mathfrak{q}}_{d}$ has the same first edge, say $e_{1}$. By Proposition 4.11, for every $d\geq 1$ we have

### v2

@@ 5.2. The property ($\ast_{k}$)
Let $k\geq 1$. A group action on a tree $(G,T)$ has property $(\ast_{k})$ if, for every geodesic $(e_{1},\ldots,e_{l+k})$ in $T$ with $l\geq 1$, we have
@@
One checks that property ($\ast_{k}$) implies property ($\ast_{k+1}$) for every $k\geq 1$.
Lemma 5.1.
Let $T$ be a tree without leaves and $G\leq\mathrm{Aut}(T)$ be a subgroup with the permutation topology induced by $\mathrm{Aut}(T)$. If $(G,T)$ has property ($\ast_{k}$) for some $k\geq 1$, then $G$ is non-discrete.
Proof.
Assume that $(G,T)$ has property $(\ast_{k})$. If we prove that all vertex-stabilisers in $G$ are infinite, then [2, Lemma 2.1] yields the claim. Let $v\in VT$. Since $T$ has no leaves, there is a ray $(e_{i})_{i\in\mathbb{Z}_{\geq 1}}$ in $T$ with $o(e_{1})=v$. For every $h\in\mathbb{Z}_{\geq 1}$, set $\mathfrak{p}_{h}=(e_{i})_{1\leq i\leq hk}$ and note that $|G_{\mathfrak{p}_{h}}\cdot(e_{hk+1},\ldots,e_{hk+k})|=|G_{\mathfrak{p}_{h}}:G_{\mathfrak{p}_{h+1}}|\geq 2$. Therefore,
$|G_{v}:G_{\mathfrak{p}_{h}}|=|G_{v}:G_{\mathfrak{p}_{1}}|\cdot\prod_{i=1}^{h-1}|G_{\mathfrak{p}_{i}}:G_{\mathfrak{p}_{i+1}}|\geq 2^{h-1},\quad\forall\,h\geq 1$
and $G_{v}$ is infinite. ∎
Lemma 5.2.
Let $T$ be a tree and $\xi$ be an end of $T$. If there is $e\in ET$ such that $G_{e}\subseteq G_{\xi}$, then $(G,T)$ does not have property ($\ast_{k}$) for any $k\geq 1$.
Proof.
Let $r=(e_{i})_{i\in\mathbb{Z}_{\geq 1}}$ be a ray in $\xi$ with $e_{1}\in\{e,\bar{e}\}$. Note that such a ray exists: indeed, we can always find a representative of $\xi$ which is completely contained either in $T_{\geq e}$ or in $T_{\geq\bar{e}}$, and $r$ is obtained by extending it to a ray starting from either $e$ or $\bar{e}$.
Since $G_{e}\subseteq G_{\xi}$, for every $g\in G_{e}$ there is $i\geq 1$ such that $g\cdot e_{j}$ is an edge of $r$ for every $j\geq i$. Since $T$ is uniquely geodesic, we conclude that every $g\in G_{e}$ fixes every edge of $r$. ∎
@@
Let $(G,T)$ be a weakly locally $\infty$-transitive action on a locally finite tree. Denote by $\omega$ the standard edge weight on $\T\Gamma=G\backslash$ T, consider $N_{\mathrm{edg}}=N_{\mathrm{edg}}^{\omega}$ as in Definition 4.4, and let $k\geq 1$. Then $(G,T)$ has property ($\ast_{k}$) if and only if $N_{\mathrm{edg}}(\rho)\geq 2$ for every path $\rho$ in $\Gamma$ of length $k+1$ that can be lifted to a geodesic in $T$.
@@
Let $k\geq 1$ and $\mathfrak{p}=(e_{1},\ldots,e_{l+k})$ be a geodesic in $T$ with $l\geq 1$. Denote by $\pi\colon T\to\Gamma$ the quotient map of $(G,T)$ and let $\pi(e_{i})=a_{i}$ for every $1\leq i\leq l+k$. By Remark 3.9.1(ii), $G_{(e_{1},\ldots,e_{l})}$ acts transitively on
$\{\mathfrak{q}=(f_{1},\ldots,f_{l+k})\in\mathrm{Geod}_{T}(e_{1}\to T)\,:\,\pi(\mathfrak{q})=\pi(\mathfrak{p})\text{ and }\forall\,i\leq l,\,f_{i}=e_{i}\}.$
Hence, by Remark 4.5.1,
(5.5) $\begin{split}&|G_{(e_{1},\ldots,e_{l})}\cdot(e_{l+1},\ldots,e_{l+k})|=\\ &=|G_{(e_{1},\ldots,e_{l})}\cdot\mathfrak{p}|=\\ &=\Big|\Big\{\mathfrak{q}=(f_{1},\ldots,f_{l+k})\in\mathrm{Geod}_{T}(e_{1}\to T):\pi(\mathfrak{q})=\pi(\mathfrak{p});\forall\,i\leq l,\,f_{i}=e_{i}\Big\}\Big|\\ &=\prod_{i=l}^{l+k-1}N_{\mathrm{edg}}(a_{i},a_{i+1})=N_{\mathrm{edg}}(a_{l},\ldots,a_{l+k}).\end{split}$
This yields the “if” part of the statement. For the “only if” part, let $\rho=(a_{1},\ldots,a_{k+1})$ be an arbitrary path in $\Gamma$ that can be lifted to a geodesic $\mathfrak{p}=(e_{1},\ldots,e_{k+1})$ in $T$. Arguing as in (5.5), we conclude that $|G_{e_{1}}\cdot\mathfrak{p}|=N_{\mathrm{edg}}(\rho)\geq 2$. ∎
@@
Let $(G,T)$ be a weakly locally $\infty$-transitive group action on a locally finite tree with quotient graph $\Gamma$ and standard edge weight $\omega$. Assume that $\omega(E\Gamma)\subseteq\mathbb{Z}_{\geq 2}$. Then the following are equivalent:
@@
Moreover, $(G,T)$ has property $(\ast_{1})$ if and only if $\omega(a)\geq 3$ for every $a\in E\Gamma$.
@@
Given $a,b\in E\Gamma$ with $t(a)=o(b)$, note that
(5.6) $N_{\mathrm{edg}}(a,b)=\omega(b)-\mathbbm{1}_{\{\bar{a}\}}(b)=1\,\Longleftrightarrow\,b=\bar{a}\text{ and }\omega(b)=2.$
[…]
@@
Let $k\geq 1$ and consider a geodesic $\mathfrak{p}=(e_{1},\ldots,e_{l+k})$ in $T$ with $l\geq 1$. By Lemma 3.1 and Corollary 4.13, there is $g\in G_{o(e_{l})}$ such that
$\begin{split}|G_{(e_{1},\ldots,e_{l})}\cdot(e_{l+1},\ldots,e_{l+k})|&=|G_{e_{l}}\cdot(e_{l+1},\ldots,e_{l+k})|\\ &=\mathcal{W}(\mathcal{L}(g\cdot(e_{l},\ldots,e_{l+k}))).\end{split}$
Since $\mathcal{L}(g\cdot(e_{l},\ldots,e_{l+k}))$ is a reduced path in $(\Delta,\iota)$ (cf. Lemma 3.5), this yields the “if” part of the statement. For the “only if” part, let $\xi=(x_{1},\ldots,x_{k+1})\in\mathcal{P}_{(\Delta,\iota)}$ and $e\in ET^{+}$ such that $\mathcal{L}(e)=x_{1}$. By Lemma 3.5, there is a geodesic $\mathfrak{p}=(e_{1},\ldots,e_{k+1})$ in $T$ with $e_{1}=e$ such that $\mathcal{L}(\mathfrak{p})=\xi$. By Proposition 4.11 we conclude that
$\mathcal{W}(\xi)=|G_{e_{1}}\cdot(e_{2},\ldots,e_{k})|\geq 2.\qed$
@@ 5.3. Convergence properties
The main goal of what follows is to study the double-coset property and the polynomial double-coset growth of triples $(G,G_{t_{1}},G_{t_{2}})$, where $G$ is a group acting on a tree $T$ as in Convention 3.0.1 and $t_{1},t_{2}\in T$.
@@
Consider a group action on a tree $(G,T)$.
@@
    Let $C:=\sup_{e,f\in ET\,:\,t(e)=o(f)}|G_{e}\cdot f|$ be finite. Then, for every geodesic $\mathfrak{p}=(e_{1},\ldots,e_{l})$ in $T$ of length $l\geq 1$, we have $|G_{e_{1}}:G_{\mathfrak{p}}|\leq C^{l-1}$.
@@
    Let $(G,T)$ have property $(\ast_{k})$ for some $k\geq 1$. Then, for every geodesic $\mathfrak{p}=(e_{1},\ldots,e_{l})$ in $T$ of length $l\geq 1$, we have $|G_{e_{1}}:G_{\mathfrak{p}}|\geq 2^{\frac{l-k}{k}}$.
@@
Since $|G_{(e_{1},\ldots,e_{i})}\cdot e_{i+1}|\leq|G_{e_{i}}\cdot e_{i+1}|$ for every $1\leq i\leq l-1$, we obtain (i).
@@
To prove the latter inequality in (5.8), set $A_{j}=|G_{(e_{1},\ldots,e_{j})}\cdot e_{j+1}|$ for every $1\leq j\leq l-1$. Then the product on the right-hand side of (5.8) becomes
@@
where, for every $1\leq j\leq l-1$,
$\alpha_{j}=|\{i\in\mathbb{Z}_{\geq 0}\,:\,\max\{1,j-k+1\}\leq i\leq\min\{j,l-k\}\}|.$
It remains to show that $\alpha_{j}\leq k$ for every $1\leq j\leq l-1$. If $j\leq k-1$, then
$\alpha_{j}=|\{i\in\mathbb{Z}_{\geq 0}\,:\,1\leq i\leq\min\{j,l-k\}\}|\leq j<k.$
If $k\leq j\leq l-k$, then $\alpha_{j}=|\{i\,:\,j-k+1\leq i\leq j\}|=k$. Finally, if $j\geq k$ and $j\geq l-k$ then $\alpha_{j}=|\{i\,:\,j-k+1\leq j\leq l-k\}|=l-j\leq k$. Hence (5.8) holds. By (5.8), the orbit-stabiliser theorem and the fact that $(G,T)$ has property $(\ast_{k})$, we conclude that
@@
Let $(G,T)$ be a group action on a locally finite leafless tree with finite quotient graph, and set $M:=\sup_{v\in VT}|o^{-1}(v)|$. If $(G,T)$ has property $(\ast_{k})$ for some $k\geq 1$, then $3\leq M<\infty$ and, for all $t_{1},t_{2}\in T$ and $n\geq 1$,
@@
In particular, for all $t_{1},t_{2}\in T$, the triple $(G,G_{t_{1}},G_{t_{2}})$ has polynomial double-coset growth.
@@
[…]
@@
$\begin{split}b_{n}(G,G_{t_{1}},G_{t_{2}})&=|\{\mathfrak{p}\in\mathrm{Geod}_{T}(T_{1}\to T)\,:\,|G_{t_{1}}:G_{\mathfrak{p}}|=n\}|\\ &\leq|\{\mathfrak{p}\in\mathrm{Geod}_{T}(T_{1}\to T)\,:\,\ell(\mathfrak{p})\leq\lfloor k\cdot\log_{2}n\rfloor+k\}|\\ &=\sum_{t\in T_{1}}\hskip-2.84544pt|\{\mathfrak{p}\in\mathrm{Geod}_{T}(t\to T)\,:\,\ell(\mathfrak{p})\leq\lfloor k\cdot\log_{2}n\rfloor+k\}|\\ &\leq\sum_{t\in T_{1}}\Bigg(1+\sum_{l=1}^{\lfloor k\cdot\log_{2}n\rfloor+k}M(M-1)^{l-1}\Bigg)\\ &=|T_{1}|\cdot\frac{M(M-1)^{\lfloor k\cdot\log_{2}n\rfloor+k}-2}{M-2}.\end{split}$
@@
Let $(G,T)$ be a weakly locally $\infty$-transitive group action on a locally finite tree with finite quotient graph $\Gamma$. Assume that the standard edge weight $\omega$ on $\Gamma$ takes values in $\mathbb{Z}_{\geq 2}$. Then the following conditions are equivalent for all $t_{1},t_{2}\in T$:
@@
    for every $a\in E\Gamma$, one has $\omega(a)\geq 3$ or $\omega(\bar{a})\geq 3$;
  • (v)
    for every $e\in ET$ and every end $\xi$ of $T$, one has $G_{e}\not\subseteq G_{\xi}$.
@@
[…]
Assume that there is $a\in E\Gamma$ such that $\omega(a)=\omega(\bar{a})=2$, and let $\mathfrak{p}=(a,\bar{a})$. Then $N_{\mathrm{edg}}(\mathfrak{p}^{d})=(\omega(\bar{a})-1)^{d}(\omega(a)-1)^{d-1}=1$ for every $d\geq 1$. In the following, $\pi\colon T\to\Gamma$ denotes the quotient map.
We first prove $\lnot$(v). Let $e\in ET$ with $\pi(e)=a$. By Remark 4.5.2, for every $d\geq 1$ there is a unique geodesic $\tilde{\mathfrak{p}}_{d}$ starting from $e$ and with $\pi(\tilde{\mathfrak{p}}_{d})=\mathfrak{p}^{d}$. By uniqueness, for all $d^{\prime}<d$, the geodesic $\tilde{\mathfrak{p}}_{d^{\prime}}$ determines the first $2d^{\prime}$ edges of $\tilde{\mathfrak{p}}_{d}$. Since $\omega(E\Gamma)\subseteq\mathbb{Z}_{\geq 2}$, there is a ray $r$ whose first $2d$ edges are the ones of $\tilde{\mathfrak{p}}_{d}$, for all $d\geq 1$. Since $N_{\mathrm{edg}}(\mathfrak{p}^{d})=1$, $G_{e}$ fixes $\tilde{\mathfrak{p}}_{d}$ for every $d\geq 1$. Hence, $G_{e}$ fixes $r$ pointwise. As a ray in $T$ belongs to the end $\xi$ precisely when it eventually equals $r$, we have $G_{e}\subseteq G_{\xi}$.
To prove $\lnot$(i), consider two paths $\mathfrak{q}_{1}=(a_{1},\ldots,a_{h})$ and $\mathfrak{q}_{2}=(b_{1},\ldots,b_{k})$ of positive length in $\Gamma$ from $\pi(t_{1})$ to $o(a)$ and from $o(a)$ to $\pi(t_{2})$, respectively. For every $d\geq 1$ we have
@@
Let $(G,T)$ be a (P)-closed group action on a locally finite tree. Assume that the quotient graph is finite and its standard edge weight takes values in $\mathbb{Z}_{\geq 2}$. Then the following conditions are equivalent for all $t_{1},t_{2}\in T$:
@@
    $(G,T)$ has property $(\ast_{k})$ for some $k\geq 1$;
  • (iv)
    for every $e\in ET$ and every end $\xi$ in $T$, one has $G_{e}\not\subseteq G_{\xi}$.
@@
[…]
Assume that $(G,T)$ does not have property $(\ast_{k})$ for every $k\geq 1$. Let $\Delta$ be the local action diagram associated to $(G,T)$ and consider an inversion $\iota$ in $\Delta$. By Proposition 5.5, there is a reduced path $(x_{1},\ldots,x_{k})$ in $(\Delta,\iota)$ of length $k\geq|X|^{2}+2$ such that $\mathcal{W}(x_{1},\ldots,x_{k})=1$, i.e., $\mathcal{W}(x_{i},x_{i+1})=1$ for every $1\leq i\leq k-1$. Since $k-1\geq|X|^{2}+1$, we have $(x_{i},x_{i+1})=(x_{i+l},x_{i+1+l})$ for some $i,l\geq 1$. Set $\eta:=(x_{j})_{i\leq j\leq i+l-1}$. For every $d\geq 1$ the power $\eta^{d}$ is a reduced path in $(\Delta,\iota)$ satisfying
@@
We first prove $\lnot$(iv). Without loss of generality (cf. Remark 5.0.1), $G=U(\Delta,\iota,c_{0})$ and $T=T(\Delta,\iota,c_{0})$, where $c_{0}\in V\Gamma$ is such that $x_{1}\in X_{c_{0}}$. Then $e=((x_{j})_{1\leq j\leq i-1},(x_{j})_{1\leq j\leq i})\in ET^{+}$. For every $d\geq 1$, denote by $\tilde{\mathfrak{p}}_{d}$ the unique geodesic in $T$ starting at $e$ and such that $\mathcal{L}(\tilde{\mathfrak{p}}_{d})=\eta^{d}$ (it exists by Lemma 3.5). By uniqueness, for every $d\geq 2$, $\tilde{\mathfrak{p}}_{d}$ prolongs $\tilde{\mathfrak{p}}_{d-1}$. Since $\mathcal{W}(\eta^{d})=1$ and the edges of $\tilde{\mathfrak{p}}_{d}$ belong to $ET^{+}$, $G_{e}=G_{\tilde{\mathfrak{p}}_{d}}$ for every $d\geq 1$ (cf. Proposition 4.11) and there is an end $\xi$ of $T$ such that $G_{e}\subseteq G_{\xi}$.
To prove $\lnot$(i), without loss of generality (cf. Remark 5.0.1), we may assume that $G=U(\Delta,\iota,c_{0})$, $T=T(\Delta,\iota,c_{0})$, either $t_{1}$ is the root of $T$ (if $t_{1}\in VT$) or the origin of $t_{1}$ is the root of $T$ (if $t_{1}\in ET$), and $\mathrm{Geod}_{T}(t_{1}\to G\cdot t_{2})\neq\emptyset$.
[…]
@@
By Lemma 3.5, for every $d\geq 1$ there is a unique geodesic $\tilde{\mathfrak{q}}_{d}\in\mathrm{Geod}_{T}(t_{1}\to G\cdot t_{2})$ with $\mathcal{L}(\tilde{\mathfrak{q}}_{d})=\xi_{1}\cdot\eta^{d}\cdot\xi_{2}$ and such that all the $\tilde{\mathfrak{q}}_{d}$’s have a common first edge, say $e_{1}\in ET^{+}$. By Remark 3.4.3, all edges of every $\tilde{\mathfrak{q}}_{d}$ are in $ET^{+}$. By Proposition 4.11, for every $d\geq 1$ we have